# Quantum Buffer Design Using Petri Nets

### v1

@@ Abstract
This paper introduces a simplified quantum Petri net (QPN) model and uses this model to generalize classical SISO, SIMO, MISO, MIMO and priority buffers to their quantum counterparts. It provides a primitive storage element, namely a quantum S-R flip-flop design using quantum CNOT and SWAP gates that can be replicated to obtain a quantum register for any given number of qubits. The aforementioned quantum buffers are then obtained using the simplified QPN model and quantum registers. The quantum S-R flip-flop and quantum buffer designs have been tested using OpenQASM and Qiskit on IBM quantum computers and simulators and the results validate the presented quantum S-R flip-flop and buffer designs.
@@ 1 Introduction
Quantum computing and information processing research has evolved over the last four decades as a viable and persistent field of interest for computer scientists, engineers, physicists and even some mathematicians[1, 2, 3, 4, 5]. What makes quantum computing and information processing attractive is the promise of quantum mechanics to provide nearly unlimited amount of parallelism in the smallest scales of physics that classical physical systems fail to offer without replicating computational resources. It is true that quantum parallelism is not suitable to handle every complex computational task but there is a sufficient set of computational problems, including factorization of large numbers, ordinary and constrained search and optimization problems, where quantum parallelism can be put to use to obtain solutions without using an exponentially increasing number of processing resources with the size of the problem in question [6, 7, 8]. There are many subfields of interest that are actively pursued in quantum computing and information processing research[9, 10, 11, 12, 13].
@@
This paper focuses on the design of quantum buffers in which classical packets (bits) are replaced by quantum packets (qubits). Buffers play a fundamental role as temporary storage elements for efficient processing of information and to synchronize the information flow between various parts of classical computer and communications systems. Consequently, their design, implementation, and performance analysis have been extensively researched in the literature[14, 15, 16, 17]. As the paradigm of computing shifts from the classical to quantum computing, conventional buffering concepts should be transformed into quantum buffering in order to pave the way for the design of future quantum computer and communication systems. To proceed in this direction, one first needs a solid model to describe the behavior of classical buffering and Petri nets largely fulfill this need in both theoretical and practical terms [18, 19, 20]. A classical Petri net is an abstract model that describes the operations of asynchronous computational structures, where tokens in places initiate (fire) transitions to form other tokens when the transition requirements are met[18]. Such requirements are usually characterized by the number of tokens entering a transition from neighboring places. The operations of various types of buffers have been modeled using Petri nets to characterize their boundedness, liveness, safeness, and stability properties and to design reliable data processing systems[21]. Recently, classical Petri nets have been extended to quantum Petri nets in which classical tokens are replaced by quantum tokens and classical transitions are replaced by quantum transitions using quantum gates[22, 23, 24]. The main objective of our work is to simplify this quantum Petri net model and use the simplified model to design quantum buffers. Quantum buffers have only been investigated in physical layers so far using fiber delay line methods[25]. Our approach is concerned with a more theoretical and design aspect of quantum buffers, where we focus on the design of quantum flip-flops, registers and buffers, rather than implementing the storage of qubits using fiber spooling and other similar delay line techniques. Our quantum buffer designs can be viewed as quantum switching networks with buffering. In this setting, quantum packets are transferred from one buffer to another using quantum tokens that hold quantum information and quantum circuits that serve as transitions that are fired when they receive the required quantum tokens. Our quantum buffer designs can also be used together with error correction schemes to build quantum circuits that are more resilient against qubit errors by protecting quantum states against the decoherence of qubits. Overall, we expect that quantum buffers will serve as temporary storage in quantum computers just like ordinary buffers in classical computers.
@@
The rest of the paper is organized as follows. The next section presents out quantum Petri net (QPN) model and provides a quantum S-R flip-flop design that can be used to build quantum registers of a desirable size. Section 3 describes the designs of SISO, SIMO, MISO, MIMO and priority quantum buffers. Section 4 provides examples of quantum buffer designs, and the paper is concluded in Section 5.
2 Quantum Petri Net Model
@@
Definition: A Quantum Petri net $(QPN)$ is a 6-tuple $(A,P,T,E,\mu,v)$ where:
@@
    $E$ is a finite set of directed and labeled arcs that connect places in $P$ with transitions in $T,$ where labels denote the $q$-tokens and their quantities for a transition to fire or to consume,
@@
The $q$-tokens in quantum Petri nets represent a set of quantum bits (qubits) or multi-qubits. Qubits are the basic units of quantum information. The $q$-tokens are represented by small black-filled circles in QPN diagrams as shown in Figure 1, where they are identified by letters $a,b,c,d,e.$ Places are locations, where $q$-tokens reside. They are represented by the large empty circles in the figure. Each place may hold a certain number of $q$-tokens, and the state of a place is determined by the $q$-tokens it contains. Places and $q$-tokens collectively represent the state of a QPN. Transitions represent quantum events (operations) that can change the state of a QPN, and manipulate $q$-tokens using quantum gates [27]. Directed arcs in $E$ link places and transitions together. The direction of an arc determines the flow of $q$-tokens. An arc from a place to a transition indicates that tokens in that place can be consumed by that transition when it fires. An arc from a transition to a place indicates that tokens can be added to that place when the transition fires. The labels on the incoming arcs to a transition such as $x$ and $y$ denote variables for $q$-tokens that are fed into a transition in the indicated quantities as seen in Figure 1. Similarly the labels on the outgoing arcs from a transition such as $z$ denote variables for $q$-tokens that are added into a place in the indicated quantities. For example, one $q$-token in each of places $P_{1}$ and $P_{2}$ ($x=a$ and $y=d$) are used by transition $T_{1}$ in Figure 1, which in this case is a controlled-not (C-NOT) gate. It is assumed that $q$-tokens are consumed by transitions in some predetermined order. In the examples presented here, an alphabetical order will be used to name tokens unless otherwise stated. When fired, a transition manipulates one or more $q$-tokens according to the function that is specified, and moves the generated $q$-tokens to the output places that are connected to that transition. For example, in Figure 1, $z=(a,a\oplus d)$ is placed in $P_{3},$ i.e., $f$ becomes $z$ after $T_{1}$ fires. Transitions in a QPN may represent quantum gates, measurements, or other quantum operations.
Figure 1: A QPN example with places $P_{1},P_{2},P_{3},$ transition $T_{1},$ and $q$-tokens $a,b,c,d,e.$
@@
Fired transitions are the only means by which $q$-tokens (qubits) move along the arcs. In quantum computing context, these transitions of $q$-tokens in a QPN represent the transformation of quantum states and how they evolve due to quantum parallelism. Effectively, extending the classical Petri net model to a quantum Petri net model adds another layer of control over quantum circuits, one in which quantum operations can be made conditional on the quantities of $q$-tokens.
@@
A marking in a QPN defines the distribution of tokens across the places that represent various configurations or states of the quantum system. The initial state of a QPN is specified by its marking that assigns $q$-tokens to its places. The evolution of the system is then represented by a sequence of changes in the states of the places, driven by the firing of transitions. As transitions fire and tokens move, they alter the overall quantum state of the system. The function $v(P_{i},\mu,t)$ represents an assignment of qubits to $q$-tokens under the marking $\mu$ at time $t$ in place $P_{i}$. Subsequent changes evolve with every transition executed, resulting in $v(P_{i},\mu,0),v(P_{i},\mu,1),\ldots$
@@
The new state of the QPN is depicted in Figure 2. In the rest of the paper, we will employ the QPN model to design four different quantum buffers.
@@
3 Quantum Buffers
For the purposes of this paper, a quantum buffer is a quantum system in which collections of qubits are structured into quantum packets, which are moved from a set of inputs to a set of outputs. Unlike classical buffers that store and process collections of bits (0’s and 1’s), quantum buffers must obey the properties of superposition, entanglement and coherence of quantum systems. The physical designs of quantum memories using various physical models have been described in [28]. Here, we focus on the conceptual design of quantum buffers using a primitive building block, called a quantum S-R (Q-S-R) flip-flop. The Q-S-R flip-flop is similar to an S-R flip-flop except that it stores a qubit rather than a bit.
Figure 3: A Quantum S-R (Q-S-R) Flip-Flop Circuit.
@@ 3 Quantum Buffers
Table 1: Behavior of a Q-S-R flip-flop.
A possible design of a Q-S-R flip-flop is depicted in Figure 3, where $|0\rangle$ and $|1\rangle$ are used as ancillary qubits to clear and set the $Q$ output. This quantum S-R flip-flop has been designed using two C-NOT and four controlled quantum swap gates with two control inputs $S$ and $R.$ The first two controlled swap gates swap their inputs when $S=|0\!\!>$ and $R=|1\!\!>,$ whereas the last two controlled swap gates swap their inputs when $S=|1\!\!>$ and $R=|0\!\!>\!\!.$ The behavior of the Q-S-R flip-flop is described in Table 1. It can be verified by tracing the qubits from the inputs to the outputs of the diagram in Figure 3. The Q-S-R flip-flops can be put together to obtain quantum registers of any desirable size as in the design of classical registers. For example, a $u$-qubit register can be put together using $u$ Q-S-R flip-flops as shown in Figure 4.
Quantum registers are building blocks of quantum buffers that store and process collections of qubits
@@
using quantum gates and circuits. The QPN model that is introduced in the earlier section facilitates this quantum processing, where quantum registers correspond to the places in the QPN model. Places can hold data $q$-tokens (multi-qubits) or ancillary $q$-tokens (multi-qubits). The ancillary $q$-tokens may represent single qubits in some buffer designs and multi-qubits in other buffer designs. They are used used to limit the capacity of a quantum buffer or select a particular transition to fire. By convention, when they represent single qubits, they will arbitrarily be initialized to $|0\!\!>$ qubit. The places that exclusively hold data $q$-tokens will be denoted by $P_{D},$ those that exclusively hold ancillary $q$-tokens will be denoted by $P_{A},$ and those that hold both data ancillary $q$-tokens will be denoted by $P_{DA}$ with appropriate indexing, such as $P_{DA,1},P_{DA,2},$ etc. The places from which $q$-tokens are fed into and taken out of the buffer will be denoted by $P_{I}$ and $P_{O},$ respectively with appropriate indexing as well.
@@ 3.1 Single Input/Single Output Quantum Buffer
Figure 5: A single-input-single-output (SISO) quantum buffer.
@@
Here, the variables with superscript * denote the qubits or multi qubits that are assigned to the data $q$-tokens, $d_{1},d_{2},\cdots,d_{n},$ at time $t=0.$ When the input transition $T_{1}$ fires, it forwards the current data $q$-token and ancillary qubit, say $d_{k}$ and $z_{k}$ to places $P_{O}$ and $P_{A1}$ respectively i.e., $f_{1}$ becomes $d_{k}$ and $f_{2}$ becomes $z_{k}$. The behavior of the SISO buffer of capacity $m=3$ is illustrated in Figure 6. The values that are listed underneath the place labels denote the number of $q$-tokens that are currently located in corresponding places. As seen at the top of Figure 6, $P_{I},P_{A},P_{A1},P_{O}$ initially hold 4, 3, 0, and 0 $q$-tokens, and the number of $q$-tokens in each place changes as the transitions $T_{1}$ continue to fire. The quantities of $q$-tokens at the end of each transition can be verified by tracing the $q$-tokens as they move between places. To process $m$ data $q$-tokens, $2m$ steps (transitions) are required and the whole process is reversible, and can be repeated again.

### v2

@@ Abstract
This paper introduces a simplified quantum Petri net (QPN) model and uses this model to generalize classical SISO, SIMO, MISO, MIMO and Priority buffers to their quantum counterparts. It provides a primitive storage element, namely a quantum S-R flip-flop and describes two different such flip-flop designs using quantum NOT, CNOT, CCNOT and SWAP gates. Each of the quantum S-R flip-flops can be replicated to obtain a quantum register for any given number of qubits. The aforementioned quantum buffers are then obtained using the simplified QPN model and quantum registers. The quantum S-R flip-flop and quantum buffer designs have been tested using OpenQasm 2.0 and Qiskit programs on IBM quantum computers and simulators and the results validate their expected operations.
@@ 1 Introduction
Quantum computing and information processing research has evolved over the last four decades as a viable and persistent field of interest for computer scientists, engineers, physicists and even some mathematicians[1, 2, 3, 4, 5]. What makes quantum computing and information processing attractive is the promise of quantum mechanics to provide nearly unlimited amount of parallelism in the smallest scales of physics that classical physical systems fail to offer without replicating computational resources. It is true that quantum parallelism is not suitable to handle every complex computational task but there is a sufficient set of computational problems, including factorization of large numbers, ordinary and constrained search and optimization problems, where quantum parallelism can be put to use to obtain solutions without using an exponentially increasing number of processing resources with the size of the problem in question [6, 7, 8]. There are several subfields of interest that are actively pursued in quantum computing and information processing research[9, 10, 11, 12, 13].
@@
This paper focuses on the design of quantum buffers in which classical packets (bits) are replaced by quantum packets (qubits). Buffers play a fundamental role as temporary storage elements for efficient processing of information and to synchronize the information flow between various parts of classical computer and communications systems. Consequently, their design, implementation, and performance analysis have been extensively researched in the literature[14, 15, 16, 17]. As the paradigm of computing shifts from the classical to quantum computing, conventional buffering concepts should be transformed into quantum buffering concepts in order to pave the way for the design of future quantum computer and communication systems. Such an effort is timely as experimental quantum memory devices have been introduced recently and suggest that reliable quantum memories will likely be available within the decade (See [18] and the extensive list of research articles cited there). To explore the utility of such devices in quantum buffering, one first needs a solid model to describe the behavior of classical buffering and Petri nets fulfill this need in both theoretical and practical terms [19, 20, 21]. A classical Petri net is an abstract model that describes the operations of asynchronous computational systems, where tokens in places initiate (fire) transitions to form other tokens when the transition requirements are met[19]. Such requirements are usually characterized by the number of tokens entering a transition from neighboring places. The operations of various types of buffers have been modeled using Petri nets to characterize their boundedness, liveness, safeness, and stability properties and to design reliable data processing systems[22]. Recently, classical Petri nets have been extended to quantum Petri nets in which classical tokens are replaced by quantum tokens and classical transitions are replaced by quantum transitions using quantum gates[23, 24, 25]. The main objective of our work is to simplify this quantum Petri net model and use the simplified model to design quantum buffers. Quantum buffers have only been investigated in physical layers so far using fiber delay line methods[26]. Our approach is concerned with a more theoretical and design aspect of quantum buffers, where we focus on the design of quantum flip-flops, registers and buffers, rather than implementing the storage of qubits using fiber spooling and other similar delay line techniques. Our quantum buffer designs can be viewed as quantum switching networks with buffering. In this setting, quantum packets are transferred from one buffer to another using quantum tokens that hold quantum information and quantum circuits that serve as transitions that are fired when they receive the required quantum tokens. Overall, we expect that quantum buffers will serve as temporary storage spaces in quantum computers just like ordinary buffers in classical computers.
@@
The rest of the paper is organized as follows. The next section presents out quantum Petri net (QPN) model and provides two quantum S-R flip-flop designs that can be used to build quantum registers of a desirable size. Section 3 describes the designs of two quantum S-R flip-flops and Section 4 presents quantum SISO, SIMO, MISO, MIMO and priority buffers. Section 5 provides examples and validations of our quantum buffer designs on IBM quantum computers and simulators. The paper is concluded in Section 6.
2 The Quantum Petri Net Model
@@
Definition: A Quantum Petri net $(QPN)$ is a 6-tuple $(D,P,T,E,\mu,v)$ where:
@@
    $E$ is a finite set of directed and labeled arcs that connect places in $P$ with transitions in $T,$ where labels denote the $q$-tokens and their quantities for a transition to consume and fire or a place to acquire,
@@
The $q$-tokens in quantum Petri nets represent a set of quantum bits (qubits) or multi-qubits. Qubits are the basic units of quantum information. The $q$-tokens are represented by small black-filled circles in QPN diagrams as shown in Figure 1, where they are identified by letters $a,b,c,d,e,$ etc. Places are locations, where $q$-tokens reside. They are represented by the large hollow circles in the figure. Each place may hold a certain number of $q$-tokens, and the state of a place is determined by the $q$-tokens it contains. Places and $q$-tokens collectively represent the overall state of a QPN. Transitions represent quantum events (operations) that can change the state of a QPN, and manipulate $q$-tokens using quantum gates [28]. Directed arcs in $E$ link places and transitions together. The direction of an arc determines the flow of $q$-tokens. An arc from a place to a transition indicates that tokens in that place can be consumed by that transition when it fires. An arc from a transition to a place indicates that tokens can be added to that place when the transition fires. As seen in Figure 1(a), multiple $q$-tokens can be added to a place at once. Figure 1(b) shows a more compact representation of the QPN in Figure 1(a), where the output arcs from transition $T_{1}$ are fused together to a single output arc with a label of $(z,2)$, indicating that the arc carries two tokens and $z=(z_{1},z_{2}).$ Henceforth, we will use the compact QPN representation. The labels on the incoming arcs to a transition such as $x$ and $y$ denote variables for $q$-tokens that are fed into a transition in the indicated quantities. Similarly the labels on the outgoing arcs from a transition such as $z$ denote variables for $q$-tokens that are added into a place in the indicated quantities. For example, one $q$-token in each of places $P_{1}$ and $P_{2}$ ($x=a$ and $y=d$) are used by transition $T_{1}$ in Figure 1, which in this case is a controlled-not (C-NOT) gate. It is assumed that $q$-tokens are consumed by transitions in some predetermined order. In the examples presented here, an alphabetical order will be used to name tokens unless otherwise stated. When fired, a transition manipulates one or more $q$-tokens according to the function that is specified, and moves the generated $q$-tokens to the output places that are connected to that transition. For example, in Figure 1, $z=(a,a\oplus d)$ is placed in $P_{3},$ i.e., $f$ becomes $z$ after $T_{1}$ fires. Transitions in a QPN may represent quantum gates, measurements, or other quantum operations.
Figure 1: (a) A QPN example with multiple arcs (b) A QPN example with compact representation of multiple arcs.
@@
Fired transitions are the only means by which $q$-tokens (qubits) move along the arcs. In quantum computing context, these transitions of $q$-tokens in a QPN represent the transformation of quantum states and how they evolve due to quantum parallelism. Effectively, extending the classical Petri net model to a quantum Petri net model adds another layer of control over quantum circuits, one in which quantum operations can be made conditional on the quantities and content of $q$-tokens.
@@
A marking in a QPN defines the distribution of tokens across the places that represent various configurations or states of the quantum system. The initial state of a QPN is specified by its marking that assigns $q$-tokens to its places. The evolution of the system is then represented by a sequence of changes in the states of the places, driven by the firing of transitions. As transitions fire and tokens move, they alter the overall quantum state of the system. The function $v(P_{i},\mu,t)$ represents an assignment of qubits to $q$-tokens under the marking $\mu$ at time $t$ in place $P_{i}$. Subsequent changes evolve with every transition executed, resulting in $v(P_{i},\mu,0),v(P_{i},\mu,1),\ldots$ The following example illustrate these ideas.
@@
The new state of the QPN is depicted in Figure 2 with the $q$-tokens $a$ and $d$ moved into place $P_{3}$ with $d$ flipped after the transition. In the rest of the paper, we will employ the QPN model to design four different quantum buffers.
@@
3 Quantum Flip-Flops
For the purposes of this paper, a quantum buffer is a quantum system in which collections of qubits are structured into quantum packets, which are moved from a set of inputs to a set of outputs. Unlike classical buffers that store and process collections of bits (0’s and 1’s), quantum buffers must obey the properties of superposition, entanglement and coherence of quantum systems. The physical designs of quantum memories using various physical models have been described in [29]. Here, we focus on the conceptual design of quantum buffers using a primitive building block, called a quantum S-R (Q-S-R) flip-flop. Quantum flip-flops are similar to classical flip-flops except that they store quantum bits rather than a classical bit. We note that quantum D, T, and J-K flip-flops have been reported earlier in the literature [30, 31] using three quantum circuits, called the R-CQCA, PPM, and F2G gates. Both R-CQCA and PPM circuits have four inputs and four outputs and can be used to create a number of quantum gates as well as design quantum flip-flops. The F2G gate has three inputs and three outputs and is used in two different designs of D, T, and J-K flip-flops. We refer the reader to [30, 31] for a detailed description of these quantum flip-flop designs. We will compare the Q-S-R flip-flops described in this paper with the two quantum J-K flip-flop designs described in  [30, 31] in Section 5 as a Q-S-R flip-flop can be obtained from a Q-J-K flip-flop rather easily by limiting the J and K inputs to $|00\rangle,|01\rangle,$ and $|10\rangle$ inputs.
Figure 3: A Quantum S-R (Q-S-R) Flip-Flop Circuit 1.
Figure 4: A Quantum S-R (Q-S-R) Flip-Flop Circuit 2.
Our designs of a Q-S-R flip-flop are depicted in Figure 3 and Figure 4 , where $|0\rangle$ and $|1\rangle$ are used as ancillary qubits to set and reset the $Q$ output. The Q-S-R flip-flop in Figure 3 has been designed using two C-NOT, one CC-NOT and four controlled quantum swap gates with two control inputs $S$ and $R.$ The first two controlled swap gates swap their inputs when $S=|0\!\!>$ and $R=|1\!\!>,$ whereas the last two controlled swap gates swap their inputs when $S=|1\!\!>$ and $R=|0\!\!>\!\!.$ The second Q-S-R flip-flop has been designed using one C-NOT, three CC-NOT gates and eight quantum NOT gates. The next state behavior of a Q-S-R flip-flop is described in Table 1. By tracing the qubits from the inputs to the outputs, it can be verified that Circuit 1 and Circuit 2 both behave like a Q-S-R flip-flop as defined in Table 1 when their $S$ and $R$ inputs are set to $|00\rangle,|01\rangle,|10\rangle.$ On the other hand, if $S$ and $R$ are set to $|11\rangle$ in the Q-S-R flip-flop in Figure 3, then the outputs $Q$ and $Q^{\prime}$ will both be forced to $|1\rangle$ if $Q=|0\rangle$ and to $|0\rangle$ if $Q=|1\rangle$, violating the assumption that $Q$ and $Q^{\prime}$ are complements of each other. More importantly, if the $S$ and $R$ inputs return to 0 about the same time then the outputs $Q$ and $Q^{\prime}$ may race to qubit $|1\rangle$ that may then potentially lead to an undefined state as in a classical S-R flip-flop. Similarly, in the Q-S-R flip-flop in Figure 4, if $S$ and $R$ are $|11\rangle$ then both outputs $Q$ and $Q^{\prime}$ will be forced to $|0\rangle$, contradicting the assumption that $Q$ and $Q^{\prime}$ oppose each other. As in the earlier case, when $S$ and $R$ inputs both go to $|0\rangle$ about the same time, the outputs $Q$ and $Q^{\prime}$ may race to qubit $|1\rangle.$
@@ 3 Quantum Flip-Flops
Table 1: The next state behavior of a Q-S-R flip-flop.
@@
4 Quantum Buffers
Quantum registers of a desirable size can be obtained using Q-S-R flip-flops as in the design of classical registers. For example, a $u$-qubit register can be put together using $u$ Q-S-R flip-flops as shown in Figure 5. Quantum registers are building blocks of quantum buffers that store and process collections of qubits using quantum gates and circuits. The QPN model that is introduced in Section 2 facilitates this quantum processing, where quantum registers correspond to the places in the QPN model. Places can hold data $q$-tokens (multi-qubits) or ancillary $q$-tokens (multi-qubits). The ancillary $q$-tokens may represent single qubits in some buffer designs and multi-qubits in other buffer designs. They are used used to limit the capacity of a quantum buffer or select a particular transition to fire. The places that exclusively hold data $q$-tokens will be denoted by $P_{D},$ those that exclusively hold ancillary $q$-tokens will be denoted by $P_{A},$ and those that hold both data ancillary $q$-tokens will be denoted by $P_{DA}$ with appropriate indexing, such as $P_{DA,1},P_{DA,2},$ etc. The places from which $q$-tokens are fed into and taken out of the buffer will be denoted by $P_{I}$ and $P_{O},$ respectively with appropriate indexing as well.
@@ 4.1 Single Input/Single Output Quantum Buffer
Figure 6: A single-input-single-output (SISO) quantum buffer. (Note: The buffer has a single port to receive data $q$-tokens from place $P_{1}$ and a single port to output them to place $P_{o}.$ The ancillary $q$-tokens are there to limit the number of transfers $P_{I}$ to $P_{o}$ to $m$ $q$-tokens as the I-gate stops firing when all $m$ ancillary qubits are consumed.)
@@
Here, the variables with superscript * denote the qubits or multi qubits that are assigned to the data $q$-tokens, $d_{1},d_{2},\cdots,d_{n},$ at time $t=0.$ When the input transition $T_{1}$ fires, it forwards the current data $q$-token and ancillary qubit, say $d_{k}$ and $z_{k}$ to places $P_{O}$ and $P_{A1}$ respectively i.e., $f_{1}$ becomes $d_{k}$ and $f_{2}$ becomes $z_{k}$. The behavior of the SISO buffer of capacity $m=3$ is illustrated in Figure 7. The values that are listed underneath the place labels denote the number of $q$-tokens that are currently located in corresponding places. As seen at the top of Figure 7, $P_{I},P_{A},P_{A1},P_{O}$ initially hold 4, 3, 0, and 0 $q$-tokens, and the number of $q$-tokens in each place changes as the transition $T_{1}$ continues to fire. The quantities of $q$-tokens at the end of each transition can be verified by tracing the $q$-tokens as they move between places. To process $m$ data $q$-tokens, $m$ steps (transitions) are required and the whole process is reversible, and can be repeated again.